trains a and b are traveling in the same direction on parallel tracks. Train a is traveling at 60mph and train b is traveling at 64mph. train a passes a station at 1:20am. If train b passes the station at 1:50am, at what time will train b catch up to train a.
Answer by Nate(3500)

Train A has a gain of 30 minutes.
D = RT
D = 60 * 0.5 = 30 miles ahead
Train A's Distance = 60t + 30
Train B's Distance = 64t
60t + 30 = 64t
30 = 4t
7.5 hours = t
7 hours and 30 minutes